### v1

@@ 1 Introduction
Large vision-language models (LVLMs) (Li et al., 2024a; Chen et al., 2024d) have achieved remarkable performance in multimodal understanding by effectively integrating visual information and user instructions into the input space of large language models (LLMs) (Grattafiori et al., 2024; Yang et al., 2025a). With the growing demand for understanding long-context visual inputs, including high-resolution images (Bai et al., 2025; Zhu et al., 2025) and long videos (Shu et al., 2025; Qin et al., 2025), LVLMs must process an increasing number of visual tokens. However, processing such long-context inputs generates numerous key-value (KV) pairs in the KV cache of LLMs, substantially increasing GPU memory consumption and degrading computational efficiency due to memory access latency and bandwidth constraints (Liu et al., 2025; Wan et al., 2024).
To address the KV storage overhead, two main approaches have emerged. Token compression methods (Yang et al., 2025b; Chen et al., 2024a) directly compress visual tokens to indirectly reduce KV cache storage, but often underperform in high-resolution fine-grained tasks like text understanding (Singh et al., 2019) and document processing (Mathew et al., 2021). More effective KV cache compression methods directly evict KV pairs in the LLM to minimize storage while preserving performance (Wan et al., 2024; Li et al., 2024b), thereby enhancing decoding efficiency and throughput. However, current KV cache compression methods for LVLMs predominantly rely on attention-based importance scores to decide which KV pairs to retain (Tao et al., 2025; Wang et al., 2025). While this strategy does reduce the KV cache size, it fails to consider the intrinsic semantic characteristics of KV pairs in multi-modal settings. To bridge this gap, we conduct a comprehensive analysis and identify two key characteristics of KV pairs in LVLMs:
@@
(II) Head-wise Redundancy Differences: Within LVLMs, different attention heads focus on distinct multi-modal aspects (Wang et al., 2025). Some heads capture global features with lower redundancy, while others focus on local details with higher semantic similarity. Figure 2 illustrates this phenomenon across multiple tasks: for Qwen2-VL-7B, certain heads show extremely high average similarity exceeding 0.9, while other heads maintain relatively low similarity below 0.3. This pattern is consistent across different vision-language tasks, indicating that KV pairs in LVLMs show varying degrees of semantic redundancy across attention heads in the LLM.
@@ 2 Related Work
Large Vision-language Models. Current large vision-language models (LVLMs) integrate a vision encoder (i.e., ViT), a projector module, and a large language model (LLM) to enable multi-modal comprehension (Liu et al., 2023; 2024a). To meet the growing demand for high-resolution image understanding, recent LVLMs introduce higher-resolution inputs via dynamic cropping strategies, such as LLaVA-NeXT (Liu et al., 2024b) and InternVL series (Chen et al., 2024c; b), or native resolution processing like Qwen2-VL (Wang et al., 2024; Bai et al., 2025) and GLM-4.5V (Hong et al., 2025). Additionally, video large language models (VideoLLMs) such as LLaVA-Video (Zhang et al., 2024) and Video-XL-2 (Qin et al., 2025) process multi-frame videos with thousands of frames. This trend dramatically increases visual token counts, leading to substantial computational costs and GPU memory burdens due to the key-value (KV) cache in the attention mechanism.
@@
KV Cache Compression. The KV cache stores computed key-value (KV) pairs during the LLM’s pre-filling phase to avoid redundant computations in decoding and enhance inference efficiency. However, long-context multi-modal inputs impose a significant GPU memory burden on the KV cache. To alleviate this, several works propose KV cache compression techniques, categorized as: (i) vision token compression that directly compresses vision tokens (Yang et al., 2025b; Chen et al., 2024a), and (ii) KV cache compression that compresses stored KV pairs during pre-filling (Zhang et al., 2023; Liu et al., 2024d). Current KV compression methods are mainly designed for LLMs, such as SnapKV (Li et al., 2024b), which clusters important KV positions using attention patterns from an observation window; KNorm (Devoto et al., 2024), which uses $\ell_{2}$ key norms to score and retain KV pairs with lower norms; and AdaKV (Feng et al., 2025), which adaptively allocates eviction budgets across attention heads. Methods specifically designed for LVLMs include InfiniPot-V (Kim et al., 2025), which employs Value Norm for KV pair selection, and SparseMM (Wang et al., 2025), which allocates asymmetric budgets across attention heads based on their importance and retains high-attention KV pairs from observation windows. Most existing compression methods follow the paradigm of retaining critical KV pairs and evicting less important ones.
@@ 3.1 Preliminary: Large Vision-language Models
KV Cache Compression. Multi-modal long-context sequences result in numerous KV pairs, which lead to significant memory overhead. KV cache compression addresses this challenge by introducing a compression operator $\bm{\Phi}$, which selectively reduces the number of stored KV cache (Wang et al., 2025). Typically, $\bm{\Phi}$ involves an evaluation function $\mathcal{E}$ that assigns scores $s_{i}=\mathcal{E}(\mathbf{K}^{l}_{h,i},\mathbf{V}^{l}_{h,i})$ to each KV pair $i$, with compression based on these scores by retaining top-$b$ highest-scoring pairs given a budget $B$, such that for each layer $l$ and head $h$, KV pairs $\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h}\in\mathbb{R}^{T\times D}$ ($T$ is sequence length, $D$ dimension per head) are compressed into compact representations $\hat{\mathbf{K}}^{l}_{h},\hat{\mathbf{V}}^{l}_{h}=\text{TopB}(\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h},\{s_{i}\}_{i=1}^{T})$. KV cache compression targets KV tensors computed during pre-filling, reducing memory burden while supporting efficient attention in decoding.
@@ 3.2 Analysis of KV Pairs Characteristics
  • Extrinsic Importance: Quantified by average attention scores from an observation window at the prompt end (default length 32) (Li et al., 2024b; Cai et al., 2025), reflecting instruction relevance. The score $s^{\text{ex}}_{\text{imp},i}=\frac{1}{|\text{window}|}\sum_{j\in\text{window}}\text{Attention}(\mathbf{Q}_{j},\mathbf{K}_{i})$ is computed from attention weights between query $\mathbf{Q}_{j}$ and key $\mathbf{K}_{i}$, balancing modalities effectively. This score serves as $s^{\text{ex}}_{\text{imp},i}$ for compression. Compression prioritizes pairs with higher $s^{\text{ex}}_{\text{imp},i}$, such that $\hat{\mathbf{K}}^{l}_{h},\hat{\mathbf{V}}^{l}_{h}=\text{TopB}(\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h},\{s^{\text{ex}}_{\text{imp},i}\}_{i=1}^{T})$, where $s^{\text{ex}}_{\text{imp},i}$ primarily reflects instruction relevance.
@@ 4.1 Experimental Setting
Model Details. We evaluate MixKV across a diverse set of architectures to ensure generalizability: LLaVA-NeXT-Mistral-7B (Liu et al., 2024b), InternVL3-8B (Li et al., 2024a), and Qwen2-VL-7B-Instruct (Wang et al., 2024) for multi-modal understanding tasks; Qwen2.5-VL-7B-Instruct (Bai et al., 2025) for GUI grounding tasks; and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) for text-only evaluation.
@@
Table 1: Performance on multiple image understanding benchmarks with reduced KV cache budgets. Since SparseMM (Wang et al., 2025) does not provide head scores for InternVL3-8B (Zhu et al., 2025), we cannot reproduce their results on this model.
@@
Implementation Details. We integrate MixKV with various KV compression methods, including SnapKV (Li et al., 2024b), PyramidKV (Cai et al., 2025), AdaKV (Feng et al., 2025), and SparseMM (Wang et al., 2025), across different KV cache budgets. Details are in Appendix A.2.
4.2 Main Results
Performance on Multi-modal Understanding Benchmarks. Table 1 presents the integration of MixKV with baseline methods on various models and benchmarks, highlighting three key advantages: (i) Universal Improvements: MixKV enhances baselines across models, benchmarks, and budgets, confirming the necessity of mixing importance with diversity. (ii) Scalability to Compression Methods: Our approach benefits various paradigms from simple baselines like SnapKV to layer-wise or head-wise budget allocation methods PyramidKV and SparseMM, modifying only the evaluation function without adjusting the compression operator. (iii) Model Compatibility: MixKV efficiently analyzes head-wise redundancy per sample, enabling direct compatibility with existing LVLMs without offline statistics like SparseMM, enabling a plug-and-play implementation.
Performance on GUI Grounding Benchmarks. Recent advancements in LVLMs demonstrate their capability to understand GUI scenarios (Cheng et al., 2024; Tang et al., 2025b; a), which rely on edge-side comprehension and necessitate KV compression to reduce storage demands. To this end, we further evaluate the fundamental GUI grounding capability on ScreenSpot-v2 using Qwen2.5-VL-7B-Instruct (Bai et al., 2025). Table 2 shows that integrating MixKV with baseline importance-based compression methods yields significant GUI grounding performance improvements. Notably, with a budget of 128, SnapKV (Li et al., 2024b) improves average precision from 75.3% to 83.3%, achieving a performance boost of +7.9%. This further validates the effectiveness of MixKV and its potential for edge-side deployment of GUI agent models.
@@ A.1 Additional Explanation of Figure 1-3
A.2 Detailed Experiment Settings
@@
  • PyramidKV (Cai et al., 2025) dynamically adjusts KV cache sizes across different LLM layers in a pyramidal manner, allocating more KV cache budget to lower layers for foundational information and less to higher layers for refined processing, based on information priority to optimize compression and performance.
@@
  • SparseMM (Wang et al., 2025) exploits sparsity patterns in visual attention heads of multi-modal models, assigning asymmetric KV cache budgets based on head importance for visual tokens, enabling modality-aware compression that effectively reduces storage requirements in vision-language models without sacrificing accuracy.
@@
A.3 More Efficiency Analysis of MixKV

### v2

@@ 1 Introduction
Large vision-language models (LVLMs) (Li et al., 2024a; Chen et al., 2024d) have achieved remarkable performance in multimodal understanding by effectively integrating visual information and user instructions into the input space of large language models (LLMs) (Grattafiori et al., 2024; Yang et al., 2025a). With the growing demand for understanding long-context visual inputs, including high-resolution images (Bai et al., 2025b; Zhu et al., 2025) and long videos (Shu et al., 2025; Qin et al., 2025), LVLMs must process an increasing number of visual tokens. However, processing such long-context inputs generates numerous key-value (KV) pairs in the KV cache of LLMs, substantially increasing GPU memory consumption and degrading computational efficiency due to memory access latency and bandwidth constraints (Liu et al., 2025c; Wan et al., 2024).
To address the KV storage overhead, two main approaches have emerged. Token compression methods (Yang et al., 2025b; Chen et al., 2024a; Liu et al., 2025b) directly compress visual tokens to indirectly reduce KV cache storage, but often underperform in high-resolution fine-grained tasks like text understanding (Singh et al., 2019) and document processing (Mathew et al., 2021). More effective KV cache compression methods directly evict KV pairs in the LLM to minimize storage while preserving performance (Wan et al., 2024; Li et al., 2024b), thereby enhancing decoding efficiency and throughput. However, current KV cache compression methods for LVLMs predominantly rely on attention-based importance scores to decide which KV pairs to retain (Tao et al., 2025; Wang et al., 2025b). While this strategy does reduce the KV cache size, it fails to consider the intrinsic semantic characteristics of KV pairs in multi-modal settings. To bridge this gap, we conduct a comprehensive analysis and identify two key characteristics of KV pairs in LVLMs:
@@
(II) Head-wise Redundancy Differences: Within LVLMs, different attention heads focus on distinct multi-modal aspects (Wang et al., 2025b). Some heads capture global features with lower redundancy, while others focus on local details with higher semantic similarity. Figure 2 illustrates this phenomenon across multiple tasks: for Qwen2-VL-7B, certain heads show extremely high average similarity exceeding 0.9, while other heads maintain relatively low similarity below 0.3. This pattern is consistent across different vision-language tasks, indicating that KV pairs in LVLMs show varying degrees of semantic redundancy across attention heads in the LLM.
@@ 2 Related Work
Large Vision-language Models. Current large vision-language models (LVLMs) integrate a vision encoder (i.e., ViT), a projector module, and a large language model (LLM) to enable multi-modal comprehension (Liu et al., 2023; 2024a). To meet the growing demand for high-resolution image understanding, recent LVLMs introduce higher-resolution inputs via dynamic cropping strategies, such as LLaVA-NeXT (Liu et al., 2024b) and InternVL series (Chen et al., 2024c; b), or native resolution processing like Qwen2-VL (Wang et al., 2024; Bai et al., 2025b) and GLM-4.5V (Hong et al., 2025). Additionally, video large language models (VideoLLMs) such as LLaVA-Video (Zhang et al., 2024b) and Video-XL-2 (Qin et al., 2025) process multi-frame videos with thousands of frames. This trend dramatically increases visual token counts, leading to substantial computational costs and GPU memory burdens due to the key-value (KV) cache in the attention mechanism.
Long-Context Optimization. Longer contexts generally improve the performance of LVLMs and enable more comprehensive multi-modal understanding (Wang et al., 2025c; Chen et al., 2025). Extensive work aims to make long-context processing more efficient and can be broadly categorized into three directions: (i) Efficient computational architectures, such as sparse attention (Li et al., 2025b; Xu et al., 2025), linear attention (Li et al., 2025a), and state-space models (Gu and Dao, 2024), which reduce the quadratic complexity of attention with respect to sequence length and thus accelerate long-context processing; (ii) Model-centric compression, including network pruning (Ma et al., 2023), model quantization (Wang et al., 2025a), and knowledge distillation (Cai et al., 2025a), which reduce parameter count and thereby lower the computational and memory cost of long-context inference; and (iii) Data-centric compression, which reduces the effective context length or storage processed by the model, for example via token compression (Yang et al., 2025b; Zhang et al., 2025), KV cache compression (Li et al., 2024b; Wang et al., 2025b) or KV cache quantization (Liu et al., 2024d; Zhang et al., 2024a), thus directly improving the efficiency of long-context computation. These three directions optimize long-context processing from complementary perspectives and are largely orthogonal to each other. Given that the context lengths required by modern applications have rapidly increased (Liu et al., 2025c), in this work we focus on the data-centric perspective and compress the stored KV cache to enable efficient long-context computation for LVLMs.
@@
KV Cache Compression. The KV cache stores computed key-value (KV) pairs during the LLM’s pre-filling phase to avoid redundant computations in decoding and enhance inference efficiency. However, long-context multi-modal inputs impose a significant GPU memory burden on the KV cache. To alleviate this, several works propose KV cache compression techniques, categorized as: (i) vision token compression that directly compresses vision tokens (Chen et al., 2024a; Liu et al., 2025a), and (ii) KV cache compression that compresses stored KV pairs during pre-filling (Zhang et al., 2023; Liu et al., 2024e). Current KV compression methods are mainly designed for LLMs, such as SnapKV (Li et al., 2024b), which clusters important KV positions using attention patterns from an observation window; KNorm (Devoto et al., 2024), which uses $\ell_{2}$ key norms to score and retain KV pairs with lower norms; and AdaKV (Feng et al., 2025), which adaptively allocates eviction budgets across attention heads. Methods specifically designed for LVLMs include InfiniPot-V (Kim et al., 2025), which employs Value Norm for KV pair selection, and SparseMM (Wang et al., 2025b), which allocates asymmetric budgets across attention heads based on their importance and retains high-attention KV pairs from observation windows. Most existing compression methods follow the paradigm of retaining critical KV pairs and evicting less important ones.
@@ 3.1 Preliminary: Large Vision-language Models
KV Cache Compression. Multi-modal long-context sequences result in numerous KV pairs, which lead to significant memory overhead. KV cache compression addresses this challenge by introducing a compression operator $\bm{\Phi}$, which selectively reduces the number of stored KV cache (Wang et al., 2025b). Typically, $\bm{\Phi}$ involves an evaluation function $\mathcal{E}$ that assigns scores $s_{i}=\mathcal{E}(\mathbf{K}^{l}_{h,i},\mathbf{V}^{l}_{h,i})$ to each KV pair $i$, with compression based on these scores by retaining top-$b$ highest-scoring pairs given a budget $B$, such that for each layer $l$ and head $h$, KV pairs $\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h}\in\mathbb{R}^{T\times D}$ ($T$ is sequence length, $D$ dimension per head) are compressed into compact representations $\hat{\mathbf{K}}^{l}_{h},\hat{\mathbf{V}}^{l}_{h}=\text{TopB}(\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h},\{s_{i}\}_{i=1}^{T})$. KV cache compression targets KV tensors computed during pre-filling, reducing memory burden while supporting efficient attention in decoding.
@@ 3.2 Analysis of KV Pairs Characteristics
  • Extrinsic Importance: Quantified by average attention scores from an observation window at the prompt end (default length 32) (Li et al., 2024b; Cai et al., 2025b), reflecting instruction relevance. The score $s^{\text{ex}}_{\text{imp},i}=\frac{1}{|\text{window}|}\sum_{j\in\text{window}}\text{Attention}(\mathbf{Q}_{j},\mathbf{K}_{i})$ is computed from attention weights between query $\mathbf{Q}_{j}$ and key $\mathbf{K}_{i}$, balancing modalities effectively. This score serves as $s^{\text{ex}}_{\text{imp},i}$ for compression. Compression prioritizes pairs with higher $s^{\text{ex}}_{\text{imp},i}$, such that $\hat{\mathbf{K}}^{l}_{h},\hat{\mathbf{V}}^{l}_{h}=\text{TopB}(\mathbf{K}^{l}_{h},\mathbf{V}^{l}_{h},\{s^{\text{ex}}_{\text{imp},i}\}_{i=1}^{T})$, where $s^{\text{ex}}_{\text{imp},i}$ primarily reflects instruction relevance.
@@ 3.3 Head-wise Adaptive Mixing Mechanism
The comprehensive score is computed through our head-wise adaptive mixing function $W^{\text{head}}(\cdot)$:
$s^{\text{comp}}_{i}=W^{\text{head}}(s_{\text{imp}}+s_{\text{div}})=(1-\bar{r}^{l}_{h})\cdot s_{\text{imp},i}+\bar{r}^{l}_{h}\cdot s^{\text{div}}_{\text{scaled},i}$ (5)
@@ 4.1 Experimental Setting
Model Details. We evaluate MixKV across a diverse set of architectures to ensure generalizability: LLaVA-NeXT-Mistral-7B (Liu et al., 2024b), InternVL3-8B (Li et al., 2024a), and Qwen2-VL-7B-Instruct (Wang et al., 2024) for multi-modal understanding tasks; Qwen2.5-VL-7B-Instruct (Bai et al., 2025b) for GUI grounding tasks; and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) for text-only evaluation.
@@
Table 1: Performance on multiple image understanding benchmarks. Since SparseMM (Wang et al., 2025b) does not provide head importance scores for InternVL3-8B (Zhu et al., 2025), we cannot reproduce their results on this model. “Full KV” means caching all KV pairs (upper bound).
@@
Implementation Details. We integrate MixKV with various KV compression methods, including SnapKV (Li et al., 2024b), PyramidKV (Cai et al., 2025b), AdaKV (Feng et al., 2025), and SparseMM (Wang et al., 2025b), across different KV cache budgets. Details are in Appendix A.3.
4.2 Main Results
Performance on Multi-modal Understanding Benchmarks. Table 1 summarizes MixKV integrated with baseline methods across models and benchmarks, and highlights three key advantages: (i) Universal improvements: MixKV consistently enhances baselines across architectures, tasks, and budgets, supporting the benefit of mixing importance with diversity. (ii) Broad applicability: it benefits a wide range of paradigms, from simple baselines such as SnapKV to layer-wise or head-wise budget allocation methods (PyramidKV and SparseMM), by modifying only the evaluation function without changing the compression operator. (iii) Model compatibility: MixKV estimates head-wise redundancy on-the-fly per sample, avoiding offline statistics (as in SparseMM) and enabling a plug-and-play integration with existing LVLMs. Additional results in Table 7 and Table 6 further verify its robustness on larger LVLMs and MoE-based models, including InternVL3-38B (Zhu et al., 2025) and Qwen3-VL-30B-A3B-Instruct (Bai et al., 2025a).
@@ 4.2 Main Results
Performance on GUI Grounding Benchmarks. Recent advancements in LVLMs demonstrate their capability to understand GUI scenarios (Cheng et al., 2024; Tang et al., 2025b; a), which rely on edge-side comprehension and necessitate KV compression to reduce storage demands. To this end, we further evaluate the fundamental GUI grounding capability on ScreenSpot-v2 using Qwen2.5-VL-7B-Instruct (Bai et al., 2025b). Table 2 shows that integrating MixKV with baseline importance-based compression methods yields significant GUI grounding performance improvements. Notably, with a budget of 128, SnapKV (Li et al., 2024b) improves average precision from 75.3% to 83.3%, achieving a performance boost of +7.9%. This further validates the effectiveness of MixKV and its potential for edge-side deployment of GUI agent models.
@@ A.1 Additional Explanation of Figure 1-3
A.2 More Discussions on Redundancy Differences
To further validate the two types of “redundancy differences” introduced in Section LABEL:sec:intro, Figure 5 visualizes the head-wise KV cache redundancy of Qwen2 (Yang et al., 2024) and Qwen2-VL (Wang et al., 2024) when processing pure-text and vision-language inputs. The pure-text and vision-language results are obtained by averaging over 100 samples randomly drawn from LongBench (Bai et al., 2024) and TextVQA (Singh et al., 2019), respectively. For pure-text inputs, Qwen2 and Qwen2-VL exhibit highly similar redundancy patterns across heads, suggesting that the architectural difference alone does not account for the redundancy gap we observe.
From this visualization, we obtain two key findings: (I) Vision-Language Redundancy Differences. When Qwen2-VL processes vision-language inputs, the semantic redundancy of its KV cache is substantially higher than for pure-text inputs, which is consistent with our analysis in the main text. Notably, for some heads (e.g., Layer 29, Heads 0 and 1), the redundancy on vision-language data is more than twice that on text-only data, reflecting the inherently redundant nature of visual signals, whereas textual tokens tend to be more semantically diverse. (II) Head-wise Redundancy Differences. For both pure-text and vision-language data, different heads exhibit markedly different redundancy levels, and their overall patterns are highly similar: a head that is relatively more redundant on text remains relatively more redundant on vision-language inputs. We hypothesize that this is because different heads focus on different types of information: some heads primarily attend to local patterns and therefore exhibit higher semantic redundancy, while others capture more global information and consequently show much lower redundancy.
A.3 Detailed Experiment Settings
@@
Figure 5: Visualization of KV cache redundancy of Qwen2 and Qwen2-VL on different data types. The number on each head is the average cosine similarity of the KV cache within that head.
@@
  • PyramidKV (Cai et al., 2025b) dynamically adjusts KV cache sizes across different LLM layers in a pyramidal manner, allocating more KV cache budget to lower layers for foundational information and less to higher layers for refined processing, based on information priority to optimize compression and performance.
@@
  • SparseMM (Wang et al., 2025b) exploits sparsity patterns in visual attention heads of multi-modal models, assigning asymmetric KV cache budgets based on head importance for visual tokens, enabling modality-aware compression that effectively reduces storage requirements in vision-language models without sacrificing accuracy.
@@
Table 5: Performance of integrating HeadKV into MixKV on LongBench and LooGLE benchmarks using Mistral-7B-Instruct-v0.2.
Methods Single-Doc QA Multi-Doc QA Avg. Long Dependency QA Avg.
NartvQA Qasper MF-en HotpotQA 2WikiMQA Musique Doc.QA Info. Retrieval Timeline Computation
Mistral-7B-Instruct-v0.2
Full KV 26.63 32.99 49.34 42.77 27.35 18.78 32.98 12.17 15.52 0.49 10.03 9.55
KV Cache Budget = 1024
HeadKV 25.88 31.28 50.54 40.61 27.57 18.80 32.45 11.93 14.87 0.49 9.56 9.21
+ MixKV 26.26 31.20 50.07 40.99 27.88 19.93 32.72 12.06 14.73 0.50 9.33 9.16
$\Delta_{\text{baseline}}$ +0.38 -0.08 -0.47 +0.38 +0.31 +1.13 +0.27 +0.13 -0.14 +0.01 -0.23 -0.05
KV Cache Budget = 128
HeadKV 24.34 26.60 48.55 40.69 25.97 15.34 30.25 10.48 12.72 0.53 10.04 8.44
+ MixKV 24.39 27.70 49.85 42.48 27.21 15.40 31.17 10.62 13.08 0.73 10.31 8.69
$\Delta_{\text{baseline}}$ +0.05 +1.10 +1.30 +1.79 +1.24 +0.06 +0.92 +0.14 +0.36 +0.20 +0.27 +0.25
A.4 Additional Experiments
Performance of MixKV with HeadKV for Long-Context Understanding. Table 5 further applies HeadKV (Fu et al., 2025) within our MixKV framework to validate its generality. Experimental results show that integrating HeadKV into MixKV can improve pure-text long-context understanding. In particular, under the extreme compression setting (i.e., Budget=128), we observe consistent and substantial gains across all tasks. This suggests that, when the KV cache budget is highly constrained, it is crucial to preserve KV entries that are both important and diverse in order to maintain the long-context understanding ability of the baseline models.
Performance of MixKV on Larger LVLMs for Multi-Modal Understanding. Table 6 reports the results of integrating MixKV with baseline KV cache compression methods on the larger LVLM InternVL3-38B (Zhu et al., 2025) to evaluate its effectiveness for multi-modal understanding. Experimental results show that MixKV consistently improves all baseline methods across benchmarks and KV cache budgets, demonstrating its robustness on larger LVLMs and further highlighting its practical value for real-world deployment.
Performance of MixKV on MoE-based LVLMs for Multi-Modal Understanding. Table 7 further presents the results of integrating MixKV with SnapKV (Li et al., 2024b) on the MoE-based LVLM Qwen3-VL-30B-A3B-Instruct to evaluate its performance on recent MoE-style LVLM architectures. Experimental results show that MixKV can significantly boost SnapKV across various benchmarks; in particular, on OCRBench, it brings a 13.6% improvement under the strict Budget=64 setting. These results demonstrate the strong effectiveness of MixKV on MoE-based LVLMs and further highlight its potential for improving the efficiency of future LVLM inference.
Table 6: Performance of Applying MixKV to InternVL3-38B.
Methods DocVQA (%) OCRBench (%) TextVQA (%) ChartQA (%) TextCaps
128 64 128 64 128 64 128 64 128 64
InternVL3-38B
Full KV 93.5 85.9 83.8 88.6 0.953
SnapKV 87.5 85.2 77.8 64.3 82.0 78.5 87.5 85.2 0.932 0.822
+ MixKV 92.1 86.9 79.3 65.8 82.8 79.4 88.2 85.8 0.959 0.859
$\Delta_{\text{baseline}}$ +4.6 +1.7 +1.5 +1.5 +0.8 +0.9 +0.7 +0.6 +0.027 +0.037
AdaKV 92.0 87.6 79.6 67.8 82.0 79.3 87.4 85.3 0.940 0.841
+ MixKV 92.3 88.5 81.1 69.2 82.9 80.2 88.2 86.0 0.961 0.859
$\Delta_{\text{baseline}}$ +0.3 +0.9 +1.5 +1.4 +0.9 +0.9 +0.8 +0.7 +0.021 +0.018
Table 7: Performance of Applying MixKV to Qwen3-VL-30B-A3B-Instruct.
Methods DocVQA (%) OCRBench (%) TextVQA (%) ChartQA (%) TextCaps
128 64 128 64 128 64 128 64 128 64
Qwen3-VL-30B-A3B-Instruct
Full KV 94.5 84.0 83.5 85.1 0.287
SnapKV 91.9 83.8 71.0 55.2 75.3 75.3 83.8 79.8 0.314 0.272
+ MixKV 93.2 86.2 80.7 68.8 80.8 79.7 84.5 80.8 0.411 0.349
$\Delta_{\text{baseline}}$ +1.3 +2.4 +9.7 +13.6 +5.5 +4.4 +0.7 +1.0 +0.097 +0.077
A.5 More Efficiency Analysis of MixKV